maggie found out that if she walks really fast to school, she can cover 2 1/2 miles in 3/4 of an hour. How fast is she walking in miles per hour?
Answer by stanbon(75887)

she can cover 2 1/2 miles in 3/4 of an hour. How fast is she walking in miles per hour?
------
(5/2)/(3/4) mph = (5/2)(4/3) mph = 10/3 mph = 3.333.. mph
-------------------------------------------------------------
Cheers,
Stan H.
---